OpenStudy (anonymous):

The area of a triangle is 96. If the base of a triangle is tripled and the height is reduced to one-third its original length, what is the new area?

OpenStudy (anonymous):

ok the base is twise the area. so it would be 192. its tripled so 192*3=576. 576 would be your base . the height would be 64 which is 1/3 of 192 which is the original base.so 577-64=513. area is half the base so 513/2=256.5 256.5 is your new area !

OpenStudy (anonymous):

My bad its 256 there is no decimal place